OpenStudy (anonymous):

divide simplify if possible.(-14/-33) / (-2/-21) type as integer or a fraction

jhonyy9 (jhonyy9):

-14/-33 * -21/-2 = -14*(-21)/(-33*(-2))= 7*7/11*1 = 49/11